OpenStudy (anonymous):

Find the surface area of a sphere with diameter 17 cm. A. 907.92 cm2 B. 226.98 cm2 C. 289 cm2 D. 106.81 cm2

OpenStudy (shamim):

u know the law of surface area of a sphere is \[A=4\pi r ^{2}\]

OpenStudy (shamim):

here A= area of ur given sphere

OpenStudy (shamim):

r= radius of ur surface

OpenStudy (shamim):

so r=17/2 cm

OpenStudy (shamim):

\[so A=4\pi \times(\frac{ 17 }{ 2 })^{2}\]

OpenStudy (shamim):

r u following me

OpenStudy (anonymous):

answer please...

OpenStudy (shamim):

a

OpenStudy (anonymous):

thank you

OpenStudy (shamim):

\[907.89cm ^{2}\]

OpenStudy (shamim):

u r most welcome

OpenStudy (shamim):

so give me a medal